OpenStudy (bobobox):

Which of the following points lie in the solution set to the following system of inequalities?(1 point) y x - 5 y -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (bobobox):

Help will medal fan and write a testimony, They put my in algebra 1 instead of pre algebra and switching me back next week, I need help to get thourgh these last assignments

OpenStudy (bobobox):

@TheSmartOne

TheSmartOne (thesmartone):

The equations are messed up from copy pasting...

OpenStudy (bobobox):

okay hold on

OpenStudy (bobobox):

@pitamar

OpenStudy (anonymous):

Back Well we have a system of 2 inequalities here: $$ y \le x - 5 \\ y \le -x - 4 $$Having a system that the variables are shared in both equations. That means that the 'y' on the top is the 'y' on the bottom and so is with the 'x'. A 'solution' to the system, is a pair of values (x,y) that when plugged in the expressions of the system they would make them true. It means that you can just take each point, plug the values into the two statements above and see if they turn out both true. For example: $$ P = (1, 10) \implies x = 1 \qquad y = 10 \\ $$Lets plug the values to the first inequality: $$ y \le x - 5 \implies 10 \le 1 - 5 \implies 10 \cancel\le -4 $$ 10 is NOT less than or equal to -4, so we get a false statement and therefore point P is not a solution for the system of equations and there is no need to go check the other.

OpenStudy (bobobox):

help me please now someone @pitamar

OpenStudy (anonymous):

Sorry, I was AFK =| @bobobox